OpenStudy (kanwal32):

If first and (2n-1)th terms of an AP,GP and H.P are equal and their n th terms are a,b,c respectively then-

OpenStudy (kanwal32):

options- A)a+2b B)a+c=b C)\[a \ge b \ge c\] D)b^2=ac

OpenStudy (kanwal32):

@ganeshie8 help @hartnn @wio @ParthKohli @paki

OpenStudy (kanwal32):

@ganeshie8 pls help

OpenStudy (anonymous):

Stop spamming chat.

OpenStudy (kanwal32):

DO U KNOW THE SOLUTION

OpenStudy (anonymous):

Yeah.

OpenStudy (kanwal32):

@wio pls help

OpenStudy (anonymous):

Do what everyone else does. Post wait, if it still isn't answered, bump it.

OpenStudy (kanwal32):

@Skrilluh I don't want ur advice

OpenStudy (anonymous):

We don't want your spam.

OpenStudy (kanwal32):

i asked @wio

OpenStudy (kanwal32):

@ganeshie8 help

OpenStudy (kanwal32):

@mathslover help

OpenStudy (kanwal32):

@wio

OpenStudy (kanwal32):

@ganeshie8

OpenStudy (anonymous):

*Sigh* if i give you the answer will you stop?

OpenStudy (kanwal32):

yeah sure

OpenStudy (anonymous):

B

OpenStudy (kanwal32):

@skillruh answer is wrong

OpenStudy (kanwal32):

@Skrilluh answer is wrong

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

You must not understand the rules of this site...

OpenStudy (kanwal32):

@ParthKohli

OpenStudy (anonymous):

Are you just listing random users?

Parth (parthkohli):

Note that the \(n\)th term of all the three series is the AM, GM and HM of the first term and the \(2n-1\)th term. Let the first and \(2n-1\)th terms be \(x,y\). It is given that:\[\dfrac{x+y}{2} = a\]\[\sqrt{xy} = b\]\[\dfrac{2xy}{x+y}=c\]From this, it is clear that \(b^2 =ac.\)